EVA Report – February 26th

Crew 263 EVA Report 26-02-2022

EVA # 6

Author: Marine Prunier

Purpose of EVA: Atmospheric Experiment Battery Change

The atmospheric experiment includes 5 devices: LOAC (Light Optical Aerosol Counter), an Electrical Field Mill, Purple Air (Aerosol Counter), Mega-Ares (Electrical Field Counter), and a Weather station. The aim of this experiment is to do testing of Martian-Atmospheric Study Equipment and their maintenance by human operators. Performing their maintenance and set-up while wearing spacesuits and following EVA protocols is the purpose of the experiment. All devices are located in the same place.

Batteries for the LOAC, the Electrical Field Mill, and Purple Air need to be changed daily. Hence, during this maintenance EVA, the 3 batteries needed to be replaced.

Start time: 08:30 am
End time: 09:00 am

Narrative: EVA Participants entered the airlock at 08:33. Depressurization ended at 08:38. EVA participants checked the rover batteries just in case. The obtained information is:

Perseverance: Usage Duration – 228.1 Battery Charge – 100%

Spirit: Usage Duration – 157.2 Battery Charge – 100%

Opportunity: Usage Duration – 84.7 Battery Charge – 100%

Curiosity: Usage Duration – 168.7 Battery Charge – 100%

The EVA participants arrived at the Atmospheric Experiment location at 08:43. The 3 batteries were replaced. Marine and Elena changed the LOAC battery. Cerise and Léa changed the Purple Air and Electrical Field Mill batteries. The status of the devices was checked.

At 08:50, the participants left the atmospheric experiment location.

At 08:55, the participants arrived at the Hab and entered Airlock.

09:00, pressurization over.

Destination: Atmospheric experience site (at the intersection between Cow Dung Road and Entrance Road)

Coordinates (use UTM NAD27 CONUS): 518288 E, 4250586 N

Participants : Marine Prunier (EVA Leader), Léa Rouverand, Cerise Cuny, Elena Lopez-Contreras Gonzalez. HabCom was Nicolas Wattelle.

Road(s) and routes per MDRS Map: Entrance Road

Mode of travel: Foot

Research Report – February 26th

Name: Marine Prunier
Crew: 263
Date: 26 Feb 2022


Robotic Telescope Requested (choose one): MLC-RCOS16

Objects to be Imaged this Evening: TOI 959.01 (Tess exoplanet catalog)

Images submitted with this report: Horsehead Nebula

Problems Encountered: /


Solar Features Observed: /

Images submitted with this report: /

Problems Encountered: /

Journalist Report – February 26th

Sol 5 – Allo… do you copy?

Author: Nicolas WATTELLE

Today we missed a rendezvous. In the context of the HF propagation study, we were supposed to contact a different “Martian base” in Toulouse. There, the radio club REF 31 should have gathered ham radios and pointed its antenna towards the MDRS.

Remember, on Sol 3 we set up the antenna near the Hab. During the following day, we set up the indoor setup: an ICOM transceiver, an antenna adapter, and a computer, linked to our triangle of steel. We did some tests, we received signals from around the planet, so we were hopeful for this morning’s session. But our experiment does not depend only on our work, another big factor named “The Sun” has also its part of the responsibility.

Mars (as Earth), has an atmosphere made of different layers. Among them, one, in particular, allows long-range propagation: the ionosphere. This layer is sensitive to Sun activity (solar storms, eruptions…), and can accumulate electrical charges. And if the total charge is significant, it permits HF waves reflections, and considerably increases the range of radiocommunications.

On Earth you might be able to perform intercontinental contacts, only using basic antennas and transceivers: a ground-ground communication without any satellite or other facility between the source and the receiver. It would be an interesting technique to use between ground stations on Mars!

So, we tried during the last months to set this experiment with the radio club (we will also perform beacon listening during the upcoming days to characterize the propagation following the propagation during the days), and today was the first attempt.

The protocol was one hour long, testing different modes of modulation, from the more reliable to the less reliable one. But the Sun might not have had enough activity the previous days: we faced a sad silence. We issued calls, but we received no answer. Have we been heard in France? We don’t know yet. But we will twice again before our mission ends, another case to follow!

That was the first experiment of the day for Valentine and me. For the rest of the crew the day was more classic: a maintenance EVA for the atmospheric experiment, adjusting the planning for the next week, repairs in the RAM…

For this one, I would like to thank a lot our Crew Engineer Matheo, who made a great job with the MegaAres antenna, tinkering with what we had in the RAM: it was out of service since the beginning of our mission, and we will be able to put it on the field by tomorrow! Thank you again Matheo (take a look at him and me, striking a pose in front of the “brand new” antenna!).

All the crew is now looking forward to tomorrow, our first “Martian rest day”! We can’t wait!

Operations Report – February 26th

Crew 263 Operations Report 26-02-2022

SOL: 5

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA


Spirit rover used: No
Hours: 157.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: handled by Shannon

Opportunity rover used: No
Hours: 84.7
Beginning charge: 100%
Ending charge: 100%
Currently charging: handled by Shannon

Curiosity rover used: No
Hours: 168.7
Beginning charge: 100%Ending charge: 100%
Currently charging: handled by Shannon

Perseverance rover used: No
Hours: 228.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: handled by Shannon

General notes and comments: NA

Summary of Hab operations:

WATER USE: 13 gallonsWater (static tank): 430 gallons
Water (loft tank): 42 gallons
Water Meter: 01561319 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No

Summary of internet: NA

Summary of suits and radios: NA

Summary of GreenHab operations: NA

Summary of ScienceDome operations:

Dual split: Off

SOC: 17% at 6:43 pm which is not normal so I notified Shannon about it.

Summary of RAM operations: Today, we used the drill, the soldering iron, and the electric fill to finish the repair of the atmospheric experiment device MegaAres that measures the local electrical field.

Summary of any observatory issues: NA

Summary of health and safety issues: See HSO Report

Questions, concerns, and requests to Mission Support: We put a trash bag in the rear airlock for pickup.

Many thanks!

Sol Summary – February 26th

Crew 263 Sol Summary Report 26-02-2022

Sol: 5

Summary Title: Back to (Martian) reality

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: In the morning we did an EVA to change the batteries of the atmospheric experiments. Nicolas and Valentine tried to make contact with Toulouse, in France (simulating a different Mars settlement) but it failed, they are going to try again in a few days.

For lunch, we cooked soup with some of the cabbage of our last harvest. It warmed up the crew members which was quite necessary as the day was very cold. Just after lunch, Shannon retrained us about EVAs and the philosophy of the sim. We were very grateful to have feedback: it will help us live the Martian experience better. We then finished the Mars puzzle together, which was quite a nice cohesion activity.

In the afternoon, we worked on our EVA requests to make them precise. Some of us went through human factors experiments. Two of the crew members used the RAM and its tools to do some repair work.

Anomalies in work: None.

Weather: Mostly sunny.

Crew Physical Status: Good.

EVA: One EVA in the morning, went well. We are doing 2 EVA Requests for tomorrow.

Reports to be filed: Sol Summary, 2 EVA Requests, EVA Report, Operations Report, Journalist Report, HSO Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site